p^2+3p-18
Can you please help me factor each trinomial competely???
I really need help with this

Answer by Alan3354(69443)

p^2+3p-18
Can you please help me factor each trinomial competely???
----------------------------------
These are almost always integers, so start looking for those. It will be:
(p + a)*(p + b), where a and b can be negative. since the coefficient of the p^2 term is 1 it's simpler.
---------------
p^2+3p-18
In this one, a*b = -18
a + b = 3
---------
The possibilities are:
1 and 18
2 and 9
3 and 6
Only 3 and 6 differ by 3, so those are the ones.
Since a*b = -18, one and not both are negative.
Since they add to +3, the larger number, 6, is positive.
So it's:
(p + 6)*(p - 3)